Bear​ Stearns' stock price closed at $96​, $101​, $57​, $33​, $3 over five successive weeks. The weekly standard deviation of the stock price calculated from this sample​ is:

The answer is 41.67 but can someone please explain this using formulas or a financial calculator. (Not the excel formula please I need to learn to do the calculations without a computer for my test)
Answer #1

average=(96+101+57+33+3)/5=58.00
Weekly Standard Deviation=(((96-58)^2+(101-58)^2+(57-58)^2+(33-58)^2+(3-58)^2)^0.5)/((5-1)^0.5)=41.67